OpenStudy (anonymous):

Identify the discontinuities, and find the zeros of each rational function. f(x) = 2x^2+5x-12/x+4

OpenStudy (anonymous):

undefined if \(x+4=0\iff x=-4\) so a discontinuity there \ zeros solve \(2x^2+5x-12=0\iff (x+4)(2x-3)=0\) so zero is at \(x=\frac{3}{2}\) but not at \(x=-4\) since the function is undefined there

OpenStudy (anonymous):

thanks